hi there! i’m stuck n solving this one using the miracle substitution(weistrass).integral of 1/(sinx(3+2cosx). i am using this as this makes everything simple. but i dont know if there are any exceptions!help me

hi there!
i’m stuck n solving this one using the miracle substitution(weistrass).integral of 1/(sinx(3+2cosx).
i am using this as this makes everything simple.
but i dont know if there are any exceptions!help me

Grade:12

1 Answers

Niranjan
8 years ago
hey take tan(x/2)=t, then sinx= 2t/(1+t^2) and cosx= (1-t^2)/(1+t^2) and dx=2dt/(1+t^2), ur integral will become ∫ 2dt/(2t*(3+2(1-t^2)))=
∫ dt/(t(5-2t^2))  . This u can solve by partial fraction...